OpenStudy (anonymous):

solve by substitution method 3x+4y=18. x=30-4y

OpenStudy (anonymous):

x=-6 y=9

OpenStudy (anonymous):

3 ( 30 - 4y) + 4y = 18

OpenStudy (anonymous):

can u show me the steps. it got to where i plug is 30-4y

OpenStudy (anonymous):

U have to substitute the x by the value 30 - 4y

OpenStudy (anonymous):

3x + 4y = 18 and x = 30 - 4y so replace the x of 3x by (30 - 4y) 3(30 - 4y) + 4y = 18 90 - 12y + 4y = 18 -8y = 18-90 y = -72/-8 y = 9

OpenStudy (anonymous):

yea ive gotten to where I sub that in. do i combine like terms next

OpenStudy (anonymous):

put y=9 in other equation

OpenStudy (anonymous):

awsome ive gotten those steps down .

OpenStudy (anonymous):

now put y = 9 in 3x + 4y = 18 3x + 4(9) = 18 3x = 18-36 x = -18/3 x = -6

OpenStudy (anonymous):

@tena i hv posted the detailed solution above.......... got a GOOD ANS

OpenStudy (anonymous):

i actually got this right lol

OpenStudy (anonymous):

good for u then....

OpenStudy (anonymous):

thanks soo much so the solution is (-6,9) or is the other way around

OpenStudy (anonymous):

thank u for the medal

OpenStudy (anonymous):

your welcome thanks soo much. my exam is stressing me out, so im trying to make sure the answers i have are correct.